# Multi-modal Generative AI: Multi-modal LLMs, Diffusions and the Unification

### v2

@@ VI-B Benchmark for the Unification
This being the case, developing such a benchmark is crucial for pushing forward the research on unification of understanding and generation, making it a promising area for future investigation.
@@ VI-C1 Leveraging multi-modal information for graph generation
Current multi-modal research predominantly focuses on modalities with regular structures with fixed degrees of freedom, e.g., texts (sequences) and images (grids). However, many real-world scenarios containing various modalities exhibit highly irregular structures with arbitrary degrees of freedom, e.g., protein structures [197], molecular graphs [198], scene graphs [199] etc. Accurately understanding and generating graphs across these modalities is an important direction for future research. For instance, Yao et al. [200] explore text-to-graph generation by leveraging the domain knowledge of LLMs, and Liu et al. [201] explore text-to-molecular graph generation by integrating the graph, image, and text information. However, there are several challenges for multi-modal graph generation: i) Understanding Structures. Given the high degree of irregularity in graphs, aligning them with various modalities poses significant difficulties. ii) Generating Structures. While mainstream approaches utilize autoregressive methods for generating discrete sequence information and employ diffusion models for generating continuous grid information, the complexity of graph structures tends to necessitate new techniques for multi-modal graph generation.
@@ VI-C2 Leveraging structural relations to facilitate multi-modal content generation
Traditional multi-modal learning methodologies often assume that data from different modalities are independent, whereas there can be strong intrinsic relationships across modalities in the real world [202, 203]. For example, the descriptions, chirps, and images of birds are more closely related to each other than those of other species such as dogs and fish. Leveraging graph structure to capture these multi-modal associations may help to understand and generate new content. Ektefaie et al. [204] explore the combination of multiple data modalities via cross-modal dependencies and geometric relationships to develop multi-modal architectures, e.g., image-intensive, knowledge-grounded, and language-intensive models, in order to process diverse datasets. Yoon et al. [205] capture intricate relationships between multiple modalities through graphs to enhance pretrained language models with multi-modal context for generative tasks. Nevertheless, several challenges remain: i) The feature spaces of different modalities are heterogeneous, thus aligning them in a unified space via a multi-modal graph poses significant challenges. ii) The connections across instances from different modalities can be heterophilous, e.g., the meow of black and white cats may be very similar, but their visual appearances differ significantly, leading to varying degrees of weights regarding similarity for the connections across modalities within the multi-modal graph. iii) There may be substantial biases among different modalities, e.g., textual and visual modalities may dominate the learning process due to the ease of collecting texts and images via the Internet, while other modalities such as acoustic perception and tactile sense are much more difficult to collect.
@@ VI-D Lightweight Multi-modal Generative AI
We define Lightweight Multi-modal Generative AI as the family of efficient Artificial Intelligence models capable of generating diverse types of data including texts, images, audios, etc., while being optimized for low computational cost, fast inference, and deployment on edge devices, e.g., smartphones, IoT devices. Lightweight Multi-modal Generative AI has broad applications in various scenarios including mobile & edge AI, IoT & embedded systems, and fast prototyping & low-cost deployment. We deem lightweight multi-modal generative AI as another promising future research direction from the following three perspectives.
1) Lightweight diffusion models face challenges from sampling steps, neural architectures, and tasks. The iterative sampling process is a critical limitation of diffusion models, bringing high computational cost and constraining real-time applications. Although substantial works (e.g., distillation [206], consistency model [207, 208] and flow matching [209, 210]) engage in few-steps (e.g., 4 steps) or single-step sampling, fewer-steps sampling in general may cause remarkable quality degradation. Tasks that require high quality [211, 212] still adopt multi-step sampling. Thus it is very important to improve few-step sampling in future investigation. Besides, the massive network architectures of diffusion models also contribute to the issue of high computational costs, which tends to be even more severe as the model size increases rapidly. Previous methods try to obtain lightweight architectures via compression techniques such as quantization [213, 214, 215], pruning [216], feature cache [217, 218], and neural architecture search [219, 220] etc. Although these works have achieved remarkable success, their designs are mostly tailored for the setting of multi-step sampling, either being not applicable or suffering from poor performances in few-steps sampling. Therefore, exploring sampling-steps-agnostic compression methods is an important future direction as well. Moreover, traditional compression methods mainly focus on UNet-based models. Existing literature [81, 78] indicates that DiT [78] may be a better architecture, resulting in the fact that more attention will be paid to DiT-based architectures. Moreover, previous compression methods mainly focus on class-condition or text-to-image generation tasks, rarely engaging in other challenging tasks such as video generation. Exploring effective compression methods for these tasks will be meaningful as well.
@@ VI-E Multi-modal Generative AI in Dynamic Environment
The multi-modal generative models discussed so far in this paper mostly do not interact with the dynamic physical world. In the future, multi-modal generative AI agents are expected to behave like humans, where they can i) perceive the multi-modal environments, ii) conduct reasoning and planning based on the perception and their current states, iii) take action to interact with the environments, and iv) improve themselves via feedbacks from the environments. A very related topic is multi-modal embodied AI [226, 227], where multi-modal LLMs are used as the controller. However, existing embodied AI methods are all parameter-fixed upon deployment, limiting their abilities of self-improvement in dynamic environments, where new concepts may arise in the course of time. The new concepts may cause the Out-of-Distribution (OOD) challenges for the pretrained multi-modal generative models, which fail to take the right action under these new concepts. Therefore, future works need to deal with the problem of i) when to update the model parameters, and ii) which part of the model parameters should be updated [228], e.g., the vision or the language modules.
@@ VII Conclusion
In this paper, we thoroughly discuss multi-modal generative AI, with a particular focus on multi-modal LLMs, multi-modal diffusion models, as well as the unifications of LLMs and diffusions for multi-modal understanding and generation. We comprehensively overview two well-documented multi-modal generative AI paradigms, i.e., multi-modal LLMs for multi-modal understanding and diffusion models for visual generation. We deeply analyze the underlying mathematical principles, fundamental architecture designs, and practical application scenarios, indicating how these models can contribute to different aspects of multi-modal generative AI. We further present the necessities for the unification of understanding and generation, exploring the theoretical possibilities and potential designs towards building unified models that jointly support understanding and generation. The unification may come across challenges such as trade-offs between autoregressive and diffusion modeling, as well as different choices between dense and MoE architectures. Beyond summarizing existing methods, we also highlight promising future directions and identify the corresponding key challenges. We believe that the discussions together with insights provided in this paper will serve as a foundation for future research and foster the development of more powerful, efficient, and generalizable multi-modal generative AI.

### v3

@@ VI-B Benchmark for the Unification
This being the case, developing such a benchmark is crucial for pushing forward the research on the unification of understanding and generation, making it a promising area for future investigation.
@@ VI-C1 Leveraging multi-modal information for graph generation
Current multi-modal research predominantly focuses on modalities with regular structures with fixed degrees of freedom, e.g., texts (sequences) and images (grids). However, many real-world scenarios containing various modalities exhibit highly irregular structures with arbitrary degrees of freedom, e.g., protein structures [193], molecular graphs [194], scene graphs [195], etc. Accurately understanding and generating graphs across these modalities is an important direction for future research. For instance, Yao et al. [196] explore text-to-graph generation by leveraging the domain knowledge of LLMs, and Liu et al. [197] explore text-to-molecular graph generation by integrating the graph, image, and text information. However, there are several challenges for multi-modal graph generation: i) Understanding Structures. Given the high degree of irregularity in graphs, aligning them with various modalities poses significant difficulties. ii) Generating Structures. While mainstream approaches utilize autoregressive methods for generating discrete sequence information and employ diffusion models for generating continuous grid information, the complexity of graph structures tends to necessitate new techniques for multi-modal graph generation.
@@ VI-C2 Leveraging structural relations to facilitate multi-modal content generation
Traditional multi-modal learning methodologies often assume that data from different modalities are independent, whereas there can be strong intrinsic relationships across modalities in the real world [198, 199]. For example, the descriptions, chirps, and images of birds are more closely related to each other than those of other species, such as dogs and fish. Leveraging graph structure to capture these multi-modal associations may help to understand and generate new content. Ektefaie et al. [200] explore the combination of multiple data modalities via cross-modal dependencies and geometric relationships to develop multi-modal architectures, e.g., image-intensive, knowledge-grounded, and language-intensive models, in order to process diverse datasets. Yoon et al. [201] capture intricate relationships between multiple modalities through graphs to enhance pretrained language models with multi-modal context for generative tasks. Nevertheless, several challenges remain: i) The feature spaces of different modalities are heterogeneous, thus aligning them in a unified space via a multi-modal graph poses significant challenges. ii) The connections across instances from different modalities can be heterophilous, e.g., the meow of black and white cats may be very similar, but their visual appearances differ significantly, leading to varying degrees of weights regarding similarity for the connections across modalities within the multi-modal graph. iii) There may be substantial biases among different modalities, e.g., textual and visual modalities may dominate the learning process due to the ease of collecting texts and images via the Internet, while other modalities, such as acoustic perception and tactile sense, are much more difficult to collect.
@@ VI-D Lightweight Multi-modal Generative AI
We define Lightweight Multi-modal Generative AI as the family of efficient Artificial Intelligence models capable of generating diverse types of data, including texts, images, audios, etc., while being optimized for low computational cost, fast inference, and deployment on edge devices, e.g., smartphones, IoT devices. Lightweight Multi-modal Generative AI has broad applications in various scenarios, including mobile & edge AI, IoT & embedded systems, and fast prototyping & low-cost deployment. We deem lightweight multi-modal generative AI as another promising future research direction from the following three perspectives.
1) Lightweight diffusion models face challenges from sampling steps, neural architectures, and tasks. The iterative sampling process is a critical limitation of diffusion models, bringing high computational cost and constraining real-time applications. Although substantial works (e.g., distillation [202], consistency model [203, 204], and flow matching [205, 85]) engage in few-steps (e.g., 4 steps) or single-step sampling, fewer-steps sampling in general may cause remarkable quality degradation. Tasks that require high quality [206, 207] still adopt multi-step sampling. Thus, it is very important to improve the few-step sampling in future investigations. Besides, the massive network architectures of diffusion models also contribute to the issue of high computational costs, which tends to be even more severe as the model size increases rapidly. Previous methods try to obtain lightweight architectures via compression techniques such as quantization [208, 209, 210], pruning [211], feature cache [212, 213], and neural architecture search [214, 215], etc. Although these works have achieved remarkable success, their designs are mostly tailored for the setting of multi-step sampling, either being not applicable or suffering from poor performances in few-step sampling. Therefore, exploring sampling-steps-agnostic compression methods is an important future direction as well. Moreover, traditional compression methods mainly focus on UNet-based models. Existing literature [98, 95] indicates that DiT [95] may be a better architecture, resulting in the fact that more attention will be paid to DiT-based architectures. Moreover, previous compression methods mainly focus on class-condition or text-to-image generation tasks, rarely engaging in other challenging tasks such as video generation. Exploring effective compression methods for these tasks will be meaningful as well.
@@ VI-E Multi-modal Generative AI in Dynamic Environment
The multi-modal generative models discussed so far in this paper mostly do not interact with the dynamic physical world. In the future, multi-modal generative AI agents are expected to behave like humans, where they can i) perceive the multi-modal environments, ii) conduct reasoning and planning based on the perception and their current states, iii) take action to interact with the environments, and iv) improve themselves via feedbacks from the environments. A very related topic is multi-modal embodied AI [221, 222], where multi-modal LLMs are used as the controller. However, existing embodied AI methods are all parameter-fixed upon deployment, limiting their abilities to self-improve in dynamic environments, where new concepts may arise in the course of time. The new concepts may cause the Out-of-Distribution (OOD) challenges for the pretrained multi-modal generative models, which fail to take the right action under these new concepts. Therefore, future works need to deal with the problem of i) when to update the model parameters, and ii) which part of the model parameters should be updated [223], e.g., the vision or the language modules.
@@ VII Conclusion
In this paper, we thoroughly discuss multi-modal generative AI, with a particular focus on multi-modal LLMs, multi-modal diffusion models, as well as the unifications of LLMs and diffusions for multi-modal understanding and generation. We comprehensively overview two well-documented multi-modal generative AI paradigms, i.e., multi-modal LLMs for multi-modal understanding and diffusion models for visual generation. We deeply analyze the underlying mathematical principles, fundamental architecture designs, and practical application scenarios, indicating how these models can contribute to different aspects of multi-modal generative AI. We further present the necessities for the unification of understanding and generation, exploring the theoretical possibilities and potential designs towards building unified models that jointly support understanding and generation. The unification may come across challenges such as trade-offs between autoregressive and diffusion modeling, as well as different choices between dense and MoE architectures. Beyond summarizing existing methods, we also highlight promising future directions and identify the corresponding key challenges. We believe that the discussions together with the insights provided in this paper will serve as a foundation for future research and foster the development of more powerful, efficient, and generalizable multi-modal generative AI.